# Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation

### v1

@@ 4 Methodology
4.1 On-Policy Learning Framework
Following previous works that applied Group Relative Proximal Optimization (GRPO) to improve the reasoning abilities of medical-VLMs, we use it to improve the reasoning of the base Med-VLM for the task of radiology report generation. GRPO [shao2024deepseekmathpushinglimitsmathematical] is an algorithm similar to Proximal Policy Optimization (PPO) [schulman2017proximal] but with two major differences: (1) Unlike PPO, it does not require a trained reward model for calculating the advantages; instead, it relies on group-relative rewards using generalized advantage estimation (GAE), and (2) it can utilize verifiable rule-based outcomes as rewards, which can be as simple as the metrics used for benchmarking and evaluation. At each alignment step, the Med-VLM $\mathcal{M}(\cdot)$ rolls out a set of candidate responses $\{o_{i}\}_{i=1}^{G}$ from the current policy $\pi_{\theta_{old}}$ and each candidate output is assigned a scalar reward $r_{i}$ at the two stages of the optimization process. The advantages $A_{i}$ are calculated using the group-relative mean and standard-deviations of the batch G,
$A_{i}\;=\;\frac{r_{i}-\operatorname{mean}\!\big(\{r_{j}\}_{j=1}^{G}\big)}{\operatorname{std}\!\big(\{r_{j}\}_{j=1}^{G}\big)}.$ (2)
Therefore, the GRPO objective is given as,
$\begin{aligned} \mathcal{L}_{\mathrm{GRPO}}(\theta)&=-\,\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\!\Big(r_{i,t}(\theta)\,\hat{A}_{i,t},\\ &\qquad\qquad\operatorname{clip}\!\big(r_{i,t}(\theta),\,1-\epsilon,\,1+\epsilon\big)\,\hat{A}_{i,t}\Big)\;-\;\beta\,D_{\mathrm{KL}}\!\left[\,\pi_{\theta}\,\middle\|\,\pi_{\mathrm{ref}}\,\right]\end{aligned}$
[…]
4.2 Supervised Fine-Tuning
@@
Before RL-tuning, we perform a supervised fine-tuning (SFT) phase to adapt $\mathcal{M}(\cdot)$ to the radiology report generation task following existing works, where SFT is carried out on curated data to improve base-reasoning and stabilize subsequent RL-based optimization [sun2024aligning, zhu2024mmedpo]. Given the SFT dataset $\mathcal{D}_{sft}=\{(I,\mathcal{R}_{\text{sft}})\}$ of chest X-rays $I$, prompts $P$, and expert reports $\mathcal{R}_{\text{sft}}$, we update the model parameters of $\mathcal{M}(\cdot)$ by minimizing the negative log-likelihood w.r.t ground-truth tokens,
$\mathcal{L}_{\mathrm{SFT}}(\theta)=\mathbb{E}_{(I,\mathcal{R}_{\text{sft}})\sim\mathcal{D}}\Bigg[-\sum_{t=1}^{T}\log\pi_{\theta}\big(r^{\text{sft}}_{t}\,\big|\,I,r^{\text{sft}}_{<t}\big)\Bigg]$
@@
where $\mathcal{R}_{\text{sft}}=(r^{\text{sft}}_{1},\dots,r^{\text{sft}}_{T})$ denotes the tokenized report, and $\pi_{\theta}$ is the conditional distribution induced by $\mathcal{M}(\cdot)$. This SFT warmup teaches the model to respect the report structure and produce structurally well-formed reports conditioned on the input image $I$ and prompt $\mathcal{P}$. In practice, we find that starting from an SFT checkpoint leads to faster convergence and avoids degenerate “cold-start” tokens. This SFT fine-tuned policy $\pi_{\theta}$ is then further trained using our multistage reward policy described below.
4.3 Multi-Stage Rewards for Visual Alignment
@@
To guide the policy toward radiology knowledge and visually grounded outputs, we design a multi-tier reward system. The first phase focuses on semantic alignment, format adherence, and clinical accuracy using textual rewards, whereas the second phase emphasizes image–text alignment between the generated report and input image, injecting disease-relevant visual reasoning.
@@
Algorithm 1 VALOR: Multi-Stage Visual Alignment
@@
Dataset $\mathcal{D}=\{(I,\mathcal{R}_{\text{gt}})\}$, Med-VLM $\mathcal{M}_{\theta}$, frozen expert $\mathcal{G}$, group size $G$ # SFT and frozen reference $\theta\leftarrow\theta_{\text{SFT}}$,  $\theta_{\text{ref}}\leftarrow\theta$,  $k\leftarrow 0$ for $t\in\{0,1\}$ do t=0t{=}0: text; t=1t{=}1: text++visual   Configure trainable modules of $\mathcal{M}_{\theta}$ for stage $t$   for minibatch $B\subset\mathcal{D}$ do    $k\leftarrow k+1$    for each $(I,\mathcal{R}_{\text{gt}})\in B$ do      Sample candidates $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid I,P)$      # Stage-1 rewards (Eq. (7))      $R_{\text{ver}}(o_{i},\mathcal{R}_{\text{gt}})\leftarrow\text{BLEU-4, ROUGE-L, BERTScore}$      $R_{\text{clin}}(o_{i},\mathcal{R}_{\text{gt}})\leftarrow s_{cx},s_{rg}$      $R_{\text{text}}(o_{i};k)\leftarrow R_{\text{ver}}+[k{=}k_{\text{clin}}]\;\lambda_{\text{clin}}R_{\text{clin}}$      if t=1t{=}1 then       # Img-text consistency (Eqn. (8)–(10))       $F_{D}\leftarrow\mathcal{G}_{\text{img}}(I)$, $e_{i}\leftarrow\mathcal{G}_{\text{txt}}(o_{i})$       $s_{i}\leftarrow\mathrm{sim}(e_{i},\mathcal{F}_{D})$, $r^{\text{vis}}_{i}\leftarrow 1-\dfrac{\mathrm{rank}(S_{i})-1}{z-1}$       $R_{\text{vis}}(o_{i};k)\leftarrow(1-\lambda_{\text{vis}})R_{\text{text}}(o_{i};k)+\lambda_{\text{vis}}\,r^{\text{vis}}_{i}$       $A_{i}\leftarrow\mathrm{RankNorm}\!\left(\{R_{\text{vis}}(o_{j};k)\}_{j=1}^{G}\right)$      else       $A_{i}\leftarrow\mathrm{RankNorm}\!\left(\{R_{\text{text}}(o_{j};k)\}_{j=1}^{G}\right)$             # GRPO policy update w. KL to reference    $\theta\leftarrow\mathrm{GRPO}\big(\theta;\{(o_{i},A_{i})\}_{i=1}^{G},\pi_{\theta_{\text{ref}}}\big)$    return $\mathcal{M}_{\theta^{\star}}$
@@
Clinically–Guided Textual Reward. In Stage-1 (t=0), we adapt the Med-VLM using verifiable text rewards that can be computed directly from the ground–truth report. We combine sentence–level BLEU-4 and ROUGE-L, which emphasize report token similarity, with BERTScore, which measures semantic similarity while being less sensitive to generic template text. In addition, we enforce a simple tag–based schema reward to discourage format violations.
@@
$\begin{aligned} R_{\text{ver}}(o_{i},g)&=\lambda_{\text{lex}}\big[\mathrm{BLEU}_{4}(o_{i},g)+\mathrm{ROUGE}_{L}(o_{i},g)\big]\\ &\quad+\lambda_{\text{sem}}\,\mathrm{BERTScore}(o_{i},g),\end{aligned}$
where $\lambda_{\text{lex}},\lambda_{\text{sem}}\!\geq 0$ and $\lambda_{\text{lex}}+\lambda_{\text{sem}}=1$.
@@
To encourage clinically correct disease terms, we use the micro-averaged F1 scores from CheXbert and RadGraph entities as periodic rewards. Let $s_{\text{cx}},s_{\text{rg}}\!\in[0,1]$ denote normalized F1 scores between labels (or triples) and entities extracted from $o$ and $g$. We define the clinical reward as,
@@
Rather than applying clinical rewards at every policy step, which can destabilize the optimization, we inject them only every $k_{\text{clin}}$ updates. Let $k$ denote the global policy update step. Then, the Stage-1 scalar reward used to compute group–relative advantages is,
@@
where $\lambda_{\text{clin}}\!\geq 0$ controls the strength of the periodic clinical contribution. This design provides dense, verifiable supervision at every step while periodically nudging the policy toward clinically faithful language without allowing sparse natural language signals to dominate training.
Visual-Alignment Reward. In Stage-2 (t=1), we introduce image-text similarity scores that explicitly tie each candidate report to the input X-Ray. For an image $I$, we obtain a continuous domain feature vector $\mathcal{F}_{D}=\mathcal{G}_{\mathrm{img}}(I)$ from the frozen expert model $\mathcal{G}$, and for each sampled candidate $o_{i}$, we compute a text embedding $\mathbf{e}_{i}^{\text{txt}}=\mathcal{G}_{\mathrm{txt}}(o_{i})$ in the same space as above. We then define the image-text similarity score for each candidate $i$,
$S_{i}\;=\;\operatorname{sim}\!\big(\mathbf{e}_{i}^{\text{txt}},\,\mathcal{F}_{D}\big),\quad i=1,\ldots,G,$ (8)
@@
To obtain a scale-free visual signal, we convert these scores into rank-normalized visual rewards. Let $\mathrm{rank}(S_{i})$ return $1$ for the largest similarity and $z$ for the smallest; we define
@@
$A_{i}^{\mathrm{vis}}\;=\;\mathrm{RankNorm}\big(\{R_{\text{vis}}(o_{j},g,I;k)\}_{j=1}^{G}\big).$ (11)
Thus, we simply replace Stage-1 advantages with $A_{i}^{\mathrm{vis}}$, which biases the policy toward candidates that are both textually faithful and visually aligned with the input image.
Format Reward. To ensure that the generated candidate reports follow the proper format, we impose a format reward (denoted as $R_{format}$) in both stages to check whether the generated report contains the required tag. Specifically, the report content must be contained within the <report> and </report> tags, respectively.
@@ 5 Experiments
5.1 Datasets
IU X-ray. Indiana University Chest X-Ray (IU-XRay) [demner2015preparing] collection contains 7470 chest X-ray images and 3955 corresponding paired reports. Each report is linked to a frontal (AP/PA) or lateral image projection. Following existing works [zhu2024mmedpo], we follow the standard 7:1:2 split.
MIMIC-CXR. MIMIC-CXR [johnson2020mimic] is one of the most extensive datasets, containing 227, 835 radiology reports from patients at the Beth Israel Deaconess Medical Center. In this work, we follow the standard convention of official train and test splits in line with the baseline methods [zhu2024mmedpo] to ensure fairness in all evaluations.
@@ 5.2 Experimental Settings
Evaluation Metrics. Following existing works [chen2020generating, chen2022cross, wang2023metransformer], we use the 14 CheXpert [smit2020chexbert] labels to assess the clinical accuracy of the generated reports. In addition, we employ commonly used natural language metrics, such as BLEU Score [papineni2002bleu], ROUGE [lin2004rouge] and METEOR [denkowski2011meteor] to assess the quality of the generated reports.
Benchmarks. We adopt the experimental settings and benchmarks as in MMedPO [zhu2024mmedpo], evaluating the accuracy and generation quality on the test splits of the IU-XRay [demner2015preparing] and MIMIC-CXR [johnson2020mimic].
Baselines and Implementation Details. We follow the same comparison protocol as in MMedPO [zhu2024mmedpo] and compare it with Direct Preference Optimization (DPO) [rafailov2023direct] a variant of preference optimization, as well as GRPO [shao2024deepseekmathpushinglimitsmathematical] without any visual alignment, as outlined in Stage-1. We use LLaVa-Medv1.5 [li2023llava] as our baseline Med-VLM, in line with other existing methods [zhu2024mmedpo]. We also compare our method on clinical accuracy against MMedPO [zhu2024mmedpo] and RADAR [hou-etal-2025-radar] using ground truth labels across 14 disease categories, evaluating precision, recall, and F-1 score. We also compare our method with SOTA proprietary models such as GPT-4o [hurst2024gpt] and MAIRA-2 for generation quality. For the GRPO fine-tuning stage, we utilize LoRA fine-tuning [hu2022lora] with a batch size of 8, a learning rate of 1e-6, and perform the alignment operation for three epochs. Further details on the benchmarking and implementation are included in the Supplementary.

### v2

@@ 4 Methodology
We describe the VALOR framework below in two parts: (1) Instruction-Following Warmup, which stabilizes the Med-VLM for structured report generation and prevents degenerate outputs during subsequent optimization, and (2) Multi-Stage Rewards for Clinical and Visual Alignment, which progressively guides the policy toward clinically precise and visually grounded reports.
4.1 Instruction-Following Warmup
@@
We perform a supervised fine-tuning (SFT) phase to enhance the instruction following capabilities of the Med-VLM $\mathcal{M}(\cdot)$ for radiology report generation. SFT is carried out on curated data to improve formatted generation and stabilize subsequent RL-based optimization [sun2024aligning, zhu2024mmedpo]. Given the SFT dataset $\mathcal{D}_{\text{sft}}=\{(\mathcal{I},\mathcal{R}_{\text{sft}})\}$ of chest X-rays $\mathcal{I}$, prompts $\mathcal{P}$, and reports $\mathcal{R}_{\text{sft}}$, we update the model parameters of $\mathcal{M}(\cdot)$ by minimizing the negative log-likelihood w.r.t. ground-truth tokens,
$\mathcal{L}_{\mathrm{SFT}}(\theta)=\mathbb{E}_{(\mathcal{I},\mathcal{R}_{\text{sft}})\sim\mathcal{D}}\Bigg[-\sum_{t=1}^{T}\log\pi_{\theta}\big(r^{\text{sft}}_{t}\,\big|\,\mathcal{I},r^{\text{sft}}_{<t}\big)\Bigg]$
@@
where $\mathcal{R}_{\text{sft}}=(r^{\text{sft}}_{1},\dots,r^{\text{sft}}_{T})$ denotes the tokenized report, and $\pi_{\theta}$ is the conditional distribution induced by $\mathcal{M}(\cdot)$. This SFT warmup teaches the model to respect the report structure and produce structurally well-formed reports conditioned on the input image $\mathcal{I}$ and prompt $\mathcal{P}$. In practice, we find that starting from an SFT checkpoint leads to faster convergence and avoids degenerate “cold-start” tokens. This SFT fine-tuned policy $\pi_{\theta}$ is then further trained using our multistage reward policy described below.
4.2 Multi-Stage Rewards for Clinical and Visual Alignment
@@
To guide the policy toward radiology knowledge and visually grounded outputs, we design a multi-tier reward system. The first stage focuses on semantic alignment, format adherence, and clinical accuracy using textual rewards, whereas the second phase emphasizes image–text alignment between the generated report and input image, injecting disease-relevant visual reasoning.
@@
Algorithm 1 VALOR: Clinically informative Multi-Stage Visual Alignment
@@
$\mathcal{D}=\{(\mathcal{I},g)\}$, Med-VLM $\mathcal{M}_{\theta}$, frozen expert $\mathcal{G}$, group size $G$ Initialize: $\theta\!\leftarrow\!\theta_{\mathrm{SFT}}$ (Eq. (2)), $\theta_{\mathrm{ref}}\!\leftarrow\!\theta$, global step $k\!\leftarrow\!0$ for $t\in\{0,1\}$ do t=0t{=}0: Stage-1 text alignment, t=1t{=}1: Stage-2 visual alignment   Configure trainable modules of $\mathcal{M}_{\theta}$ for stage $t$   for minibatch $\mathcal{B}\subset\mathcal{D}$ do    $k\leftarrow k+1$    for each $(\mathcal{I},g)\in\mathcal{B}$ do      Sample $G$ candidate reports $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid\mathcal{I},\mathcal{P})$      Compute $R_{\mathrm{ver}}(o_{i},g)$ (Eq. (3)) and $R_{\mathrm{clin}}(o_{i},g)$ (Eq. (4))      Form Stage-1 reward $R_{\mathrm{text}}(o_{i},g;k)$ using periodic clinical injection (Eq. (5))      if $t=1$ then       Compute image/text features and similarity $S_{i}$ (Eq. (6)); rank-normalize to $r_{i}^{\mathrm{vis}}$ (Eq. (7))       Compose $R_{\mathrm{vis}}(o_{i},g,\mathcal{I};k)$ (Eq. (8)) and set $A_{i}\leftarrow A_{i}^{\mathrm{vis}}$ via Eq. (9)      else       Set $A_{i}\leftarrow\mathrm{RankNorm}\!\big(\{R_{\mathrm{text}}(o_{j},g;k)\}_{j=1}^{G}\big)$ Adv. computation             Update policy with KL regularization to $\pi_{\theta_{\mathrm{ref}}}$ (Eq. (10))    return $\mathcal{M}_{\theta^{\star}}$
@@
Clinically Informed Textual Reasoning. Here, we adapt the Med-VLM using verifiable text rewards that can be computed directly from the ground truth report. We combine NLG Metrics [papineni2002bleu, lin2004rouge], which emphasize report token similarity, with completeness [zhang2019bertscore], which measures semantic coherence while being less sensitive to generic template text. At each alignment step, $\mathcal{M}(\cdot)$ rolls out a set of candidate reports $\{o_{i}\}_{i=1}^{G}$ from the current policy $\pi_{\theta}$. For each candidate report $o_{i}$ and ground truth report $g$, the verifiable text reward is given as,
@@
$\begin{aligned} R_{\text{ver}}(o_{i},g)&=\lambda_{\text{lex}}\big[\mathrm{BLEU}_{4}(o_{i},g)+\mathrm{ROUGE}_{L}(o_{i},g)]+\lambda_{\text{sem}}\,\mathrm{BERTScore}(o_{i},g),\end{aligned}$
where weight hyperparameters $\lambda_{\text{lex}},\lambda_{\text{sem}}\!\geq 0$ and $\lambda_{\text{lex}}+\lambda_{\text{sem}}=1$.
@@
To encourage clinically correct disease terms, we use the micro-averaged F1 scores from CheXbert and RadGraph entities as periodic rewards. Let $s_{\text{cx}},s_{\text{rg}}\!\in[0,1]$ denote normalized F1 scores between labels (or triplets) and entities extracted from $o_{i}$ and $g$. We define the clinical reward as,
@@
Rather than applying clinical rewards at every policy step, which destabilizes the optimization, we inject them only every $k_{\text{clin}}$ updates. Let $k$ denote each rollout from the global policy update step. Then, the Stage-1 scalar reward used to compute group relative advantages is combined as,
@@
where $\lambda_{\text{clin}}\!\geq 0$ controls the strength of the periodic clinical contribution. This design provides dense, verifiable supervision at every step while periodically nudging the policy toward clinically faithful language without allowing natural language metrics to entirely dominate the training process.
Self-Supervised Visual Reasoning. In Stage-2, we introduce image-text similarity scores that explicitly tie each candidate report with the provided X-Ray. For an image $I$, we obtain a continuous domain feature vector $\mathcal{F}_{D}=\mathcal{G}(\mathcal{I})$ from the frozen expert model, and for each sampled candidate $o_{i}$, we compute a text embedding $\mathbf{e}_{i}^{\text{txt}}=\mathcal{G}_{\mathrm{txt}}(o_{i})$ in the same space as above. We then define the image-text similarity score for each candidate $i$,
$S_{i}\;=\;\operatorname{sim}\!\big(\mathbf{e}_{i}^{\text{txt}},\,\mathcal{F}_{D}\big),\quad i=1,\ldots,G$
@@
To obtain a scale-free visual signal, we convert these scores into rank-normalized visual rewards. Let $\mathrm{rank}(S_{i})$ return $1$ for the largest similarity and $z$ for the smallest; we define,
@@
$A_{i}^{\mathrm{vis}}\;=\;\mathrm{RankNorm}\big(\{R_{\text{vis}}(o_{j},g,I;k)\}_{j=1}^{G}\big).$
Thus, we simply replace Stage-1 advantages with $A_{i}^{\mathrm{vis}}$, which biases the policy toward candidates that are both textually faithful and visually aligned. Therefore, the optimization objective is given as,
$\begin{aligned} \mathcal{L}_{\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\texttt{VALOR}}}}(\theta)&=-\,\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\!\Big(r^{vis}_{i,t}(\theta)\,A^{vis}_{i,t},\\ &\qquad\qquad\operatorname{clip}\!\big(r^{vis}_{i,t}(\theta),\,1-\epsilon,\,1+\epsilon\big)\,A^{vis}_{i,t}\Big)\;-\;\beta\,D_{\mathrm{KL}}\!\left[\,\pi_{\theta}\,\middle\|\,\pi_{\mathrm{ref}}\,\right],\end{aligned}$
where, $D_{\mathrm{KL}}\!\big(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\big)$ serves as a regularization term penalizing divergence of the current policy $\pi_{\theta}$ from the reference policy $\pi_{\mathrm{ref}}$, with $\beta$ controlling the regularization strength w.r.t. the policy.
Format Reward. To ensure that generated candidate reports follow the proper format, we impose a format reward (denoted as $R_{\text{format}}$) in both stages to check whether the generated report contains the required tag. Specifically, the report must be contained within the <report> and </report> tags, respectively.
@@ 5.1 Experimental Settings
Datasets and Evaluation Metrics. We benchmark VALOR on two widely used radiology report datasets: IU-XRay [demner2015preparing] and MIMIC-CXR [johnson2020mimic]. The generated reports are evaluated for quality across two critical categories: clinical accuracy and natural language generation quality. For clinical accuracy, we follow existing works [chen2020generating, chen2022cross, wang2023metransformer], using 14 CheXpert [smit2020chexbert] labels to assess the clinical accuracy of the generated reports. Addtionally, we also evaluate on more clinically oriented metrics such as GREEN [ostmeier2024green] and RaTEScore [zhao2024ratescore]. In addition, we employ commonly used natural language metrics, such as BLEU Score [papineni2002bleu], ROUGE [lin2004rouge] and METEOR [denkowski2011meteor] to assess the quality of the generated reports.
Figure 4: Generated reports with VALOR. First Case: MedVLM [li2023llava] fails to detect subtle reasoning patterns needed for accurate diagnosis, resulting in hallucination of disease observations such as ‘‘consolidation’’. In contrast, VALOR with verifiable guidance leads to significantly better reasoning and diagnosis capabilities. With visual alignment, VALOR’s enhanced multimodal reasoning leads to correct diagnosis of all observations, showcasing the need for awareness of disease-relevant portions and focused attention. The MedVLM again hallucinates diseases due to insufficient visual focus, in comparison VALOR reduces hallucinations and generates reports faithful to the X-Ray exhibiting better anatomical and semantic awareness. Second Case: Similarly in the second case, although MedVLM is able to deduce that the X-ray is benign, it’s reasoning is incorrect as highlighted in it’s generated report. VALOR comes to the same conclusion but demonstrating better overall reasoning capabilities.
Baselines and Implementation Details. We compare against a set of direct baselines on top of the base Med-VLM including SFT, DPO [rafailov2023direct] and GRPO [shao2024deepseekmathpushinglimitsmathematical]. We also compare against recent post-alignment work MMedPO [zhu2024mmedpo] which uses Direct Preference Optimization (DPO), as well as GRPO without any visual alignment and clinical intervention, as outlined in Stage-1. We adopt LLaVa-Medv1.5 [li2023llava] as our baseline Med-VLM, in line with [zhu2024mmedpo] and provide additional results with QwenVL-2.5-7B [bai2025qwen3] in the supplementary. We also extensively compare our method on clinically informative using ground truth labels across 14 disease categories, evaluating F1 score, GREEN [ostmeier2024green] and RaTEScore [zhao2024ratescore]. Further, we compare VALOR against zero-shot non-medical SOTA proprietary models such as GPT-4o [hurst2024gpt], GPT-4.1 [achiam2023gpt] Gemini-2.5 Pro [comanici2025gemini] and medical SOTA proprietary models MedFlamingo [moor2023med], MAIRA-2 [bannur2024maira] and HuatuoGPT-V-8B [chen2024towards] for generation quality. For the GRPO training phase, we utilize LoRA fine-tuning [hu2022lora] with a batch size of 8, a learning rate of 1e-6, and perform the alignment operation for three epochs. Further details on the benchmarking, instruction prompts and additional implementation details are included in the supplementary.